seamus is using all 120inches of a piece of copper wire to build a rectangle that is five times as long as it is wide and a square whose side lenght is the same as the width of the rectangle what is the area of the square?
Answer by TimothyLamb(4379)

L = 5W
W = (1/5)L
s = W
4s + 2L + 2W = 120
---
2W + 2L + 2W = 120
4W + 2L = 120
(4/5)L + (10/5)L = 120
(14/5)L = 120
L = 120(5/14)
L = 600/14
---
L = 300/7
s = 60/7
---
Answer:
a = 3600/49 = 73.47 sq.in.
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php